OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

OpenStudy (anonymous):

means if you replace the variable in the equation by the number, you get a true equality. for example, the number \(5\) is a solution to the following \[x=5\] \[2x=10\] \[2x+4=14\] \[2(2x+4)=28\] \[2(2x+4)-5=23\] and so on